OpenStudy (anonymous):

Need help with a lollipop problem >.< i'll post it up in a bit. Please and Thank you

OpenStudy (anonymous):

here you go

OpenStudy (anonymous):

ignore the picture on the right

OpenStudy (ranga):

V = 4/3(pi)r^3 Take derivative with respect to time t dV/dt = 4/3(pi)(3r^2)dr/dt = 4(pi)r^2dr/dt Just for visualization you can think of the sphere as made up of liquid and each lick subtract the same few drops of water. So dV/dt can be assumed to be a constant. K = 4(pi)r^2dr/dt dr/dt = K / ((4(pi)r^2) = C / r^2 The rate of decrease in radius is inversely proportional to the square of the radius. So when radius is high, dr/dt will be small. When radius is small, dr/dt will be pretty high.

OpenStudy (anonymous):

our teacher said it was constant

OpenStudy (anonymous):

@razor99

razor99 (razor99):

yep

OpenStudy (anonymous):

@razor99 the kids are right, right ?

razor99 (razor99):

yep they are

OpenStudy (anonymous):

yayy we were right :D Thanks so much \(^-^)/

OpenStudy (anonymous):

@razor99 can you explain this part dr/dt = K / ((4(pi)r^2) = C / r^2 well the C part

razor99 (razor99):

well um i am not godd at that one :(

razor99 (razor99):

*good

OpenStudy (anonymous):

did you mean k instead of C ?? for the bottom???? dr/dt = K / ((4(pi)r^2) = C / r^2

OpenStudy (anonymous):

@razor99

razor99 (razor99):

um yes

OpenStudy (anonymous):

thnx

razor99 (razor99):

no problem

OpenStudy (ranga):

dr/dt = K / (4(pi)r^2) = C / r^2 K, 4, pi are all constants and can be grouped together as another constant C. C = K / (4(pi))